# QCD predictions for physical multimeson scattering amplitudes

### v1

@@ Abstract
Existing LQCD calculations of multihadron processes have predominantly focused on two-particle decay channels and often used quark masses heavier than their physical values. By reducing the quark masses and incorporating decay modes involving more than two hadrons, it is possible to explore well-known three-pion resonances such as the $\omega(782)$ and $a_{1}(1260)$, exotic tetraquarks such as $T_{\rm cc}\to DD\pi$, the enigmatic Roper resonance, $N(1440)\to N\pi+N\pi\pi$, and even three-nucleon forces. The successful development of three-body tools for LQCD calculations [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45] has transformed these aspirations into a practical reality [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63].
@@
This Letter is released in parallel with a long article, in which all details of the calculations are presented. Here we present the highlights and a sketch of the methods.

### v2

@@ Abstract
Existing LQCD calculations of multihadron processes have predominantly focused on two-particle decay channels and often used quark masses heavier than their physical values. By reducing the quark masses and incorporating decay modes involving more than two hadrons, it is possible to explore well-known three-pion resonances such as the $\omega(782)$ and $a_{1}(1260)$, exotic tetraquarks such as $T_{\rm cc}\to DD\pi$, the enigmatic Roper resonance, $N(1440)\to N\pi+N\pi\pi$, and even three-nucleon forces. The successful development of three-body tools for LQCD calculations [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45] has transformed these aspirations into a practical reality [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64].
@@
This Letter is released in parallel with a long article with all the details of the calculations [65]. Here we present the highlights and an outline of the methods.